An investment of $2500 is made at an annual simple interest rate of 5.5%. How much additional money must be invested at an annual simple interest rate of 8% so that the total interest earned is 7% of the total investment?
Answer by TimothyLamb(4379)

x = amount invested at 5.5%
y = amount invested at 8%
---
x = 2500
0.055x + 0.08y = 0.07*(2500 + y)
0.055x + 0.08y = 0.07*2500 + 0.07y
0.055x + 0.01y = 175
---
put the system of linear equations into standard form
---
x = 2500
0.055x + 0.01y = 175
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = amount invested at 5.5% = 2500
y = amount invested at 8% = 3750
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---